### v1

# Discovering a Novel Dynamics Mechanism for Charmonium Scattering

Abstract
In this letter, we propose a novel mechanism for charmonium scattering that utilizes the internal structure of charmonium. By capturing light flavor quarks and anti-quarks from the vacuum, charm and anti-charm quarks form virtual charmed mesons, which mediate an effective one-boson exchange process. This approach accurately reproduces the di-$J/\psi$ invariant mass spectrum observed by CMS and LHCb, demonstrating its validity. Our mechanism offers a comprehensive framework for understanding charmonium scattering and is applicable to the scattering problems involving all fully heavy hadrons, an area of increasing interest.
Introduction.—Scattering problems, as a fundamental topic in physics, offering critical insights into the interactions and internal structures of matter. Using this approach, Yukawa predicted the pion meson Yukawa:1935xg , which mediates the nuclear force, establishing a cornerstone of modern nuclear physics. In this century, advances in experimental techniques have led to the discovery of numerous new hadronic states, presenting unique opportunities to explore hadron scattering. Noteworthy achievements include the identification of exotic molecular states such as $P_{c}$ states LHCb:2015yax ; LHCb:2019kea and $T_{cc}(3875)^{+}$  LHCb:2021vvq ; LHCb:2021auc . These discoveries not only mark the advent of ’Particle 2.0’ but also enhance our understanding of the nonperturbative aspects of the strong force Chen:2016qju ; Richard:2016eis ; Chen:2016spr ; Lebed:2016hpi ; Esposito:2016noz ; Guo:2017jvc ; Ali:2017jda ; Brambilla:2019esw ; Liu:2019zoy ; Chen:2022asf .
Recently, the CMS Collaboration reported a high-precision measurement of the invariant mass spectrum of di-$J/\psi$ CMS:2023owd , confirming the previous observation of $X(6900)$ by the LHCb Collaboration LHCb:2020bwg and revealing additional enhancements in the di-$J/\psi$ invariant mass spectrum. Furthermore, ATLAS has joined this effort, alongside LHCb and CMS ATLAS:2023bft . This novel finding surprised the community, prompting various explanations, including exotic hadronic states liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Chen:2020xwe ; Deng:2020iqw ; Gordillo:2020sgc ; Zhao:2020zjh ; Faustov:2021hjs ; Wang:2021kfv and coupled-channel dynamics generation Wang:2020wrp ; Wang:2020tpt ; Wang:2022jmb ; Gong:2020bmg ; Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr ; Lu:2023aal ; Guo:2020pvt ; Zhuang:2021pci . However, a substantial issue—charmonium scattering—has not been fundamentally resolved, despite some tentative efforts Wang:2020wrp ; Wang:2020tpt ; Wang:2022jmb ; Gong:2020bmg ; Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr ; Lu:2023aal ; Guo:2020pvt ; Zhuang:2021pci . The highly effective one-boson ($\pi$, $\sigma$, $\rho$, $\cdots$) exchange mechanism for describing hadron scattering fails in quantitatively characterizing charmonium scattering, as charmonium is a typical example of a fully heavy hadron. Thus, studying how charmonium undergoes scattering has become a particularly typical issue in current scattering problems and urgently requires a solution.
Figure 1: A novel approach enabling one-boson exchange for effective charmonium scattering. The cross symbol on the left indicates that direct light flavor one-boson exchange ($\pi,\rho,\sigma,~{}\cdots$) cannot occur.
In this letter, we propose a novel mechanism for charmonium scattering. As illustrated in Fig. 1, the charm and anti-charm quarks within charmonium can capture light flavor anti-quark and quark, respectively, created from the vacuum to form virtual charmed mesons. These virtual mesons provide a source that enables one-boson exchange to be effective for charmonium scattering. A direct application of this novel mechanism successfully reproduces the di-$J/\psi$ invariant mass spectrum data from CMS and LHCb, demonstrating the mechanism’s validity. We should emphasize that this mechanism is not limited to charmonium scattering; it can also be applied to investigate the scattering problems involving all fully heavy hadrons, a topic that is gaining attention from the community, including lattice research groups Lyu:2021qsh ; Mathur:2022ovu .
Application to reproduce the CMS and LHCb data.—The di-$J/\psi$ invariant mass spectrum observed by the CMS and LHCb Collaborations exhibit significant features around the dip near 6.8 GeV and the peak near 6.9 GeV, corresponding closely to the $\psi\psi(3686)$ and $\psi\psi(3770)$ thresholds, respectively. In contrast, the bumps near 6.6 GeV and 7.3 GeV do not align with any thresholds involving particles from the $J/\psi$ family. Therefore, to capture the main features, we consider couplings among the three channels: ${J/\psi J/\psi,J/\psi\psi(3686),J/\psi\psi(3770)}$.
When one-boson exchange is introduced at the quark level, as depicted in Fig. 1, the scattering mechanism between two charmonia can proceed as shown in Fig. 2. Initially, the two charmonia, $\psi_{1}$ and $\psi_{2}$, couple into pairs of virtual $D$ mesons. Within each pair, one $D$ meson acts as a spectator, while the other engages in an interaction mediated by the exchange of a light hadron, $\mathcal{P}$. Subsequently, the interacting virtual $D$ mesons combine with the spectators to form the final charmonia, $\psi_{3}$ and $\psi_{4}$.
Figure 2: The scattering mechanisms of the $\psi_{1}\psi_{2}\to\psi_{3}\psi_{4}$ process involve $D$ meson loops, with diagram $(a)$ representing the $t$-channel contribution and diagram $(b)$ representing the $u$-channel contribution. Here, $\psi_{i}$ can be $J/\psi$, $\psi(3686)$, or $\psi(3770)$, and $\mathcal{P}$ can be $\eta$ or $\sigma$. Furthermore, when $\psi_{1}=\psi_{2}$ and $\psi_{3}=\psi_{4}$, there is no $u$-channel contribution.
As shown in Fig. 2, we have neglected the contributions from charmed-strange meson loops. This is primarily due to the significantly lower creation rate of $s\bar{s}$ quark pairs compared to $u\bar{u}$ or $d\bar{d}$ quark pairs, with a ratio of approximately ${1}/{\sqrt{3}}$. Consequently, contributions from charmed-strange mesons are at least ${2}/{3}$ smaller. Additionally, the higher threshold for producing charmed-strange meson pairs leads to further suppression by the loop function. Therefore, to focus on the primary physics, we exclude contributions from charmed-strange meson loops.
Furthermore, Fig. 2 illustrates that the exchanged light hadron, $\mathcal{P}$, is limited to $\sigma$ and $\eta$. This restriction arises because, considering the loop mechanism as a black box, the interaction between two charmonia and a light hadron must comply with the conservation laws of the strong interaction. Given that the mass of the exchanged particle should be as small as possible, only $\sigma$ and $\eta$ are viable candidates.
We perform the entire calculation by adopting the effective Lagrangian approach, with all relevant Lagrangians being Cheng:1992xi ; Yan:1992gz ; Wise:1992hn ; Burdman:1992gh ; Casalbuoni:1996pg
$\mathcal{L}_{D^{(\ast)}D^{(\ast)}\eta}$ $=$ $-\frac{2g}{\sqrt{6}f_{\pi}}\left(i\varepsilon_{\alpha\mu\nu\lambda}v^{\alpha}D^{\ast\mu}D_{a}^{\ast\lambda}\partial^{\nu}\eta+DD_{\lambda}^{*}\partial^{\lambda}\eta\right),$ (1)
$\mathcal{L}_{D^{(\ast)}D^{(\ast)}\sigma}$ $=$ $-2g_{s}DD\sigma+2g_{s}D^{\ast}\cdot D^{\ast}\sigma,$ (2)
$\mathcal{L}_{D^{\ast}\bar{D}^{\ast}J/\psi}$ $=$ $-ig_{D^{\ast}D^{\ast}\psi}\left[\psi\cdot\bar{D}^{\ast}\overleftrightarrow{\partial}\cdot D^{\ast}-\psi^{\mu}\bar{D}^{\ast}\cdot\overleftrightarrow{\partial}{}^{\mu}D^{\ast}\right.$ (3)
$\left.\left.+\psi^{\mu}\bar{D}^{\ast}\cdot\overleftrightarrow{\partial}D^{\ast\mu}\right)\right],$
$\mathcal{L}_{D^{\ast}\bar{D}J/\psi}$ $=$ $g_{D^{\ast}D\psi}\epsilon_{\beta\mu\alpha\tau}\partial^{\beta}\psi^{\mu}\left(\bar{D}\overleftrightarrow{\partial}{}^{\tau}D^{\ast\alpha}+\bar{D}^{\ast\alpha}\overleftrightarrow{\partial^{\tau}}D\right),$ (4)
$\mathcal{L}_{D\bar{D}J/\psi}$ $=$ $ig_{DD}\psi\cdot\bar{D}\overleftrightarrow{\partial}D.$ (5)
Here, $v\equiv i\overleftrightarrow{\partial}/2\sqrt{m_{i}m_{f}}$, with $m_{i,f}$ representing the masses of the initial or final heavy meson Ding:2021igr . For the coupling constants $g$, $f_{\pi}$, and $g_{s}$, we follow Ref. Ding:2021igr ; Falk:1992cx ; Isola:2003fh ; Liu:2009qhy ; Chen:2019asm and set them as $g=0.59$, $f_{\pi}=132$ MeV, and $g_{s}=0.76$, respectively. The coupling constants $g_{D^{\ast}D^{\ast}\psi}$, $g_{DD\psi}$, and $g_{D^{\ast}D\psi}$, according to heavy quark effective theory, can be related to the parameter $f_{\psi}$ as $\frac{g_{D^{\ast}D^{\ast}\psi}}{m_{D^{\ast}}}=\frac{g_{DD\psi}}{m_{D}}=g_{D^{\ast}D\psi}=\frac{m_{\psi}}{m_{D}f_{\psi}}$.
The amplitudes for the diagrams shown in Fig. 2 can be generally expressed as
$\mathcal{M}_{(a)}$ $=$ $16\epsilon_{1}^{\mu}\epsilon_{2}^{\nu}\epsilon_{3}^{\ast\alpha}\epsilon_{4}^{\ast\beta}\frac{1}{q^{2}-m_{\mathcal{P}}^{2}+im_{\mathcal{P}}\Gamma_{\mathcal{P}}}\mathcal{F}^{2}_{q}(q)$ (6)
×$\times\int\frac{d^{4}q_{1}}{(2\pi)^{4}}\frac{\mathcal{V}_{\mu\alpha}}{\prod_{i}(p_{i}^{2}-m_{i}^{2}+im_{i}\Gamma_{i})}\mathcal{F}_{q_{1}}(q_{1})$
×$\times\int\frac{d^{4}q_{2}}{(2\pi)^{4}}\frac{\mathcal{V}_{\nu\beta}}{\prod_{j}(p_{j}^{2}-m_{j}^{2}+im_{j}\Gamma_{j})}\mathcal{F}_{q_{2}}(q_{2}),$
$\mathcal{M}_{(b)}$ $=$ $16\epsilon_{1}^{\mu}\epsilon_{2}^{\nu}\epsilon_{4}^{\ast\alpha}\epsilon_{3}^{\ast\beta}\frac{1}{q^{2}-m_{\mathcal{P}}^{2}+im_{\mathcal{P}}\Gamma_{\mathcal{P}}}\mathcal{F}_{q}^{2}(q)$ (7)
×$\times\int\frac{d^{4}q_{1}}{(2\pi)^{4}}\frac{\mathcal{V}_{\mu\alpha}}{\prod_{i}(p_{i}^{2}-m_{i}^{2}+im_{i}\Gamma_{i})}\mathcal{F}_{q_{1}}(q_{1})$
×$\times\int\frac{d^{4}q_{2}}{(2\pi)^{4}}\frac{\mathcal{V}_{\nu\beta}}{\prod_{j}(p_{j}^{2}-m_{j}^{2}+im_{j}\Gamma_{j})}\mathcal{F}_{q_{2}}(q_{2}),$
where $\mathcal{V}_{\gamma\delta}$ denotes the Lorentz structure of the loop integral, $\mathcal{F}_{q}(q)$ is the form factor for the exchanged light hadron $\mathcal{P}$ to account for the off-shell effect, and $\mathcal{F}_{q_{i}}(q_{i})$ is a form factor used to avoid ultraviolet divergence in the loop integral. We use the widely adopted monopole form for $\mathcal{F}_{q}(q)$, given by $\mathcal{F}_{q}(q)=\frac{\Lambda_{q}^{2}-m_{q}^{2}}{\Lambda_{q}^{2}-q^{2}}$, where $\Lambda_{q}$ is parametrized by a free parameter $\alpha_{q}$ as $\Lambda_{q}=m_{q}+\alpha_{q}\Lambda{\mathrm{QCD}}$, with $\Lambda{\mathrm{QCD}}=220$ MeV and $\alpha_{q}$ is an adjustable parameter. For $\mathcal{F}_{q_{i}}(q_{i})$, to avoid possible unphysical poles on the real axis of the momentum sheet, we adopt $\mathcal{F}_{q_{i}}(q_{i})=\exp(-\vec{q}_{i}^{~{}2}/\Lambda_{i}^{2})$.
Additionally, the reason for multiplying an overall factor of 16 in both Eq. (6) and Eq. (7) is that each $D$ meson loop in Fig. 2 can involve not only $D^{(\ast)}D^{(\ast)}\bar{D}^{(\ast)}$ configurations but also $\bar{D}^{(\ast)}\bar{D}^{(\ast)}D^{(\ast)}$ configurations. Furthermore, for both of these cases, there are charged and neutral configurations. Thus, there are a total of 4 configurations for each loop, resulting in 16 configurations for each diagram in Fig. 2, all of which have the same value.
Using Eqs. (6)-(7), we can express the effective potential in momentum space as
$\mathcal{V}_{(i)}(\sqrt{s},\vec{k},\vec{k}^{\prime})=-\frac{\mathcal{M}_{(i)}(\sqrt{s},\vec{k},\vec{k}^{\prime})}{2\sqrt{p_{1}^{0}p_{2}^{0}p_{3}^{0}p_{4}^{0}}}.$ (8)
Here, $(i)$ stands for $(a)$ or $(b)$ given in Fig. 2, $\sqrt{s}$ is the center-of-mass energy, $\vec{k}$ and $\vec{k}^{\prime}$ are the three-momenta of the initial and final particles in the center-of-mass frame as $\vec{k}=(0,0,k)$ and $\vec{k}^{\prime}=(k^{\prime}\sin\theta,0,k^{\prime}\cos\theta)$, respectively, and $p_{j}^{0}$ denotes the energy of particle $j$.
In this letter, we consider only the $S$-wave behavior in the di-$J/\psi$ spectrum. Therefore, we first perform a partial wave expansion, which can be calculated as
$\mathcal{V}^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})\sim 2\pi\int d\cos\theta d_{\lambda\lambda^{\prime}}^{J}(\theta)\mathcal{V}_{\lambda,\lambda^{\prime}}(\sqrt{s},\vec{k},\vec{k}^{\prime})$ (9)
with $k^{(\prime)}=|\vec{k}^{~{}(\prime)}|$, where $\lambda$ and $\lambda^{\prime}$ is the helicity indices of the initial and final states, respectively, and $d_{\lambda,\lambda^{\prime}}^{J}(\theta)$ is the Wigner-$d$ matrix.
Next, we substitute the obtained partial wave effective potentials into the coupled-channel Lippmann-Schwinger equation
$T^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})$ $=$ $\mathcal{V}^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})+\sum_{\lambda^{\prime\prime}}\int\frac{k^{\prime\prime 2}dk^{\prime\prime}}{(2\pi)^{3}}$ (10)
×$\times\mathcal{V}^{J}_{\lambda,\lambda^{\prime\prime}}(\sqrt{s},k,k^{\prime\prime})G_{0}(\sqrt{s},k^{\prime\prime})$
×$\times T^{J}_{\lambda^{\prime\prime},\lambda^{\prime}}(\sqrt{s},k^{\prime\prime},k^{\prime}),$
then, we can extract the on-shell coupled-channel $T$-matrices $T^{J}(\sqrt{s})$. Following a similar approach as in Refs. Gong:2020bmg ; Dong:2020nwy , the transition amplitude can be written as
$\mathcal{M}^{J}_{J/\psi J/\psi}(\sqrt{s})=\alpha e^{-\beta s}\left(1+\sum_{i}r_{i}e^{i\phi_{i}}G_{0i}(\sqrt{s})T_{i,J/\psi J/\psi}^{J}(\sqrt{s})\right),$
where $\alpha$ represents an overall factor, $r_{i}$ and $\phi_{i}$ describe the relative production strength and phase angle for channel $i$, respectively, and $\beta$ is used to simulate the background.
Finally, the di-$J/\psi$ spectrum can be calculated as Gong:2020bmg ; Dong:2020nwy
$\sigma(s)=\frac{\left|\vec{p}_{J/\psi}\right|}{8\pi s}\left|\mathcal{M}^{J}_{J/\psi J/\psi}(\sqrt{s})\right|^{2},$ (11)
with $\vec{p}_{J/\psi}$ being the three momentum of final $J/\psi$ in the center-of-mass frame.
We primarily focus on the CMS data, utilizing the LHCb data as supplementary due to the higher precision of the CMS muon detector. The fit parameters are as follows, excluding the semi-fixed overall factors $\alpha$ and $\beta$ which control the lineshape and background magnitude. We first select a set of $(\alpha,\beta)$ values, fix them, perform the fit, and then choose the best fit result for presentation. For each channel $i$, there are two parameters: the phase angle $\phi_{i}$ and the relative production strength $r_{i}$.
Our model includes cut-off dependent form factors, making the scattering amplitudes inherently cut-off dependent. To avoid lengthy fitting processes, we fix all cut-offs ($\alpha_{q}$ and $\Lambda_{i}$) to one and treat the decay constants $f_{\psi}$ as free parameters. This approach absorbs the cut-off dependence into the coupling constants, adhering to the principles of renormalization. Consequently, there are three additional parameters: $f_{J/\psi}$, $f_{\psi(3686)}$, and $f_{\psi(3770)}$, resulting in a total of 11 parameters for our fit.
For the di-$J/\psi$ invaraint mass spectrum, considering only the $S$-wave effective potential in the scattering, there are two possible quantum numbers: $0^{++}$ and $2^{++}$. The $1^{++}$ configuration is excluded because, although $1\otimes 1\rightarrow 1$ is possible, it cannot form a state with $C=+$ parity under $S$-wave conditions. Tab. 1 presents the results of the two sets of fit parameters for the $0^{++}$ and $2^{++}$ configurations. The corresponding fit results, based on the parameters in Tab. 1, are shown in Fig. 3.
Table 1: Fit parameters for reproducing the CMS experimental data in the $0^{++}$ and $2^{++}$ configurations. Parameters marked with an asterisk ($\ast$) are fixed during the fit.
Parameter $0^{++}$ $2^{++}$
$\alpha$ 617.338 981.439
$\beta$ 0.0447 0.0528
$r_{J/\psi J/\psi}$ 1.101 0.429
$\phi_{J/\psi J/\psi}$ 2.979 3.179
$r_{J/\psi\psi(3686)}$ 0.724 0.487
$\phi_{J/\psi\psi(3686)}$ 2.484 2.745
$r_{J/\psi\psi(3770)}$ 1.508 0.268
$\phi_{J/\psi\psi(3770)}$ 0.196 0.542
$f_{J/\psi}$ 0.378 GeV 0.428 GeV
$f_{\psi(3686)}$ 0.237 GeV 0.116 GeV
$f_{\psi(3770)}$ 0.760 GeV 0.267 GeV
$\chi^{2}/d.o.f$ 2.067 2.268
Figure 3: The fit results on experimental data are presented, with the top panel for $0^{++}$ configuration and the bottom panel for $2^{++}$ configuration. In each panel, the dark and light grey error bars represent the CMS data and LHCb data respectively, and the red solid line indicates the fit result. The magenta dashed line, light blue line filled with shadows, light green dot-dashed line, and orange dotted line show the contributions of background, $J/\psi J/\psi$ channel, $J/\psi\psi(3686)$ channel, and $J/\psi\psi(3770)$ channel, respectively.
As shown in Fig. 3, our model successfully reproduces the overall features of the di-$J/\psi$ invariant mass spectrum. This indicates that our proposed scattering mechanism between two charmonium states is reasonable, providing a novel approach to understanding hadron scattering, especially for fully heavy hadrons. Furthermore, both the lineshape of the black solid lines and the $\chi^{2}/d.o.f$ values in Table 1 suggest a slight preference for the $0^{++}$ configuration in the di-$J/\psi$ invariant mass spectrum. However, the $\chi^{2}/d.o.f$ values are not close to 1, primarily because we did not include channels with thresholds near 6.6 GeV and 7.3 GeV, leading to mismatches in these regions. Despite this, our primary goal of validating the new scattering mechanism is achieved through the correspondence with experimental results.
In addition, Fig. 3 reveals a significant twist at 6.865 GeV, spanning about 10 MeV, in the partial contributions of some channels, such as the $J/\psi\psi(3686)$ channel in the $0^{++}$ fit. This is attributed to an anomalous threshold effect caused by $\psi(3770)$. Since $\psi(3770)$ is above the $D\bar{D}$ threshold, it can decay into a pair of real $D$ mesons when it interacts with another vector charmonium, as illustrated in Fig. 3. When $\psi_{3}=J/\psi$ and the remaining intermediate meson in the triangle loop is also a $D$ meson, the anomalous threshold effect occurs, resulting in the observed twist around 6.9 GeV. This kind of threshold effect has been extensively discussed in previous studies Lu:2023aal ; Zhuang:2021pci , and our results are consistent with Ref. Lu:2023aal . The presence of such an effect is a crucial feature of our model. More precise future experiments around 6.9 GeV would greatly aid in validating our model.
Moreover, the wide peak around 6.6 GeV can be reproduced by interference effects without introducing additional channels. The sharp resonance-like structure in $\sigma_{J/\psi J/\psi}$ in both configurations suggests a potential $J/\psi J/\psi$ bound state, as it is not reflected in the overall spectrum due to phase space limitations. This conclusion aligns with Refs. Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr . Importantly, the dip near 6.8 GeV, particularly evident in the $J/\psi\psi(3686)$ contribution in the $2^{++}$ fit, indicates the possible existence of a state similar to $\Upsilon(10753)$, which also shows as a dip in $e^{+}e^{-}\to b\bar{b}$ cross section Dong:2020tdw . This pole around 6.8 GeV is also identified in Refs. Dong:2020nwy ; Liang:2021fzr ; Guo:2020pvt ; Zhuang:2021pci . Thus, detailed future experimental analyses of this dip could further test our model.
Overall, our model not only reproduces the experimental data but also integrates nearly all conclusions from previous studies into a unified framework. This initial success in understanding the scattering mechanism between two charmonium states motivates further study and broader application of our model. This letter represents the first step in that direction.
Summary.—Recently, the CMS Collaboration reported a high-precision measurement of the di-$J/\psi$ invariant mass spectrum CMS:2023owd , which confirmed the previous observation of $X(6900)$ by the LHCb Collaboration LHCb:2020bwg and revealed additional enhancements. The ATLAS Collaboration has also contributed to this effort, reinforcing these findings ATLAS:2023bft . These observations have spurred a variety of theoretical interpretations, including the existence of exotic hadronic states and dynamics generated by coupled channels liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Chen:2020xwe ; Deng:2020iqw ; Gordillo:2020sgc ; Zhao:2020zjh ; Faustov:2021hjs ; Wang:2021kfv . However, a critical issue remains unresolved: the nature of charmonium scattering. Traditional one-boson exchange mechanisms, highly effective for describing hadron scattering, fall short in quantitatively characterizing charmonium scattering. This shortcoming is particularly pronounced in fully heavy hadron systems.
In this letter, we introduce a novel mechanism for charmonium scattering. Our approach leverages the internal structure of charmonium, where charm and anti-charm quarks can capture light flavor anti-quarks and quarks from the vacuum, forming virtual charmed mesons. These virtual mesons act as mediators, enabling an effective one-boson exchange process for charmonium scattering.
Our proposed mechanism successfully reproduces the di-$J/\psi$ invariant mass spectrum data from CMS and LHCb CMS:2023owd ; LHCb:2020bwg , validating its effectiveness. A significant advantage of this approach is its broad applicability. It is not limited to charmonium scattering but can be extended to investigate the scattering of all fully heavy hadrons, an area of increasing interest within the physics community, including lattice QCD research groups Lyu:2021qsh ; Mathur:2022ovu . This approach addresses the limitations of traditional one-boson exchange mechanisms, which are inadequate for fully heavy hadron systems due to the absence of light quark exchanges. By incorporating virtual meson formation, our model provides a robust framework for understanding charmonium interactions.
This study provides a comprehensive framework for charmonium scattering, integrating well with existing experimental observations CMS:2023owd ; LHCb:2020bwg and theoretical efforts on scattering mechanisms Wang:2020wrp ; Wang:2020tpt ; Wang:2022jmb ; Gong:2020bmg ; Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr ; Lu:2023aal ; Guo:2020pvt ; Zhuang:2021pci . The success of our mechanism in explaining the di-$J/\psi$ spectrum sets a precedent for future studies on fully heavy hadron scattering. We emphasize the importance of further experimental investigations, particularly high-precision measurements around key energy thresholds, to refine and validate our model. Our work lays the groundwork for a deeper understanding of scattering mechanisms in fully heavy hadron systems, with potential implications for broader applications in particle physics.
Acknowledgements.—This work is supported by the National Natural Science Foundation of China under Grant Nos. 12335001, 12247101, 12305139, and 12305087, the National Key Research and Development Program of China under Contract No. 2020YFA0406400, and the 111 Project under Grant No. B20063, the Fundamental Research Funds for the Central Universities, and the project for top-notch innovative talents of Gansu province. R. C. is supported by the Xiaoxiang Scholars Programme of Hunan Normal University. Q. H. is supported by the Start-up Funds of Nanjing Normal University under Grant No. 184080H201B20.